if a football field is 40 yds wide and 120 yds long. what is the distance between opposite corners of the football field?

Answer by checkley79(3341)

YOU HAVE A RIGHT TRIANGLE WITH SIDES OF 40 & 120 YDS.
A^2+B^2=C^2
40^2+120^2=C^2
1600+14400=C^2
16000=C^2
C=126.49 YARDS ANS.